OpenStudy (waheguru):

Please Help Me Fast ! -50 - 4 ( (-10-4 (-3+1)^2 ) / (-2) )

OpenStudy (waheguru):

Can someone help me solve this step by step

OpenStudy (waheguru):

anyone

OpenStudy (anonymous):

What is -3 + 1 = ??

OpenStudy (waheguru):

-2 but I dont get how to eliminate the brackets

OpenStudy (anonymous):

It is in the bracket now.. What is the square of -2??

OpenStudy (waheguru):

-50 - 4 ( (-10 - 4(-2)^2 ) / (-2) )

OpenStudy (waheguru):

After this I am getting confused because dont we take the brackets out so it beccomes (-10-4+4)

OpenStudy (waheguru):

but the answers ahve ( -10 -4(4)

OpenStudy (waheguru):

I am just talking about the middle part

OpenStudy (anonymous):

-50-4[(-10-4(-3+1)^2] -54[-10-4(-2)^2] -54[-10-4(4)] -54(-10-16) -54(-26) 1404

OpenStudy (waheguru):

why is -54[-10-4(4)] the (4) in brackets dont they get removed

OpenStudy (waheguru):

After the operation is done do we not remove the brackets also or do they stay? no matter what happends

OpenStudy (anonymous):

\[ (-10-4 (-3+1)^2 ) / (-2) = \frac{-10-4(4)}{-2} = \frac{-26}{-2} = 13\]

OpenStudy (waheguru):

I am just confused about the brackets, do we keep them and when do we take the brackets out ?

OpenStudy (waheguru):

I dont know how to word my question but it is about the brackets and when we remove them ?

OpenStudy (anonymous):

can u right ur equation clearly?? :)

OpenStudy (eyust707):

Once the parenthesis are simplified you may remove them a replace them with multiplication.

OpenStudy (eyust707):

In other situations you may also distribute to rid the parenthesis.

OpenStudy (anonymous):

So, now the expression has become: \[\huge \color{green}{-50 -4(13) = ???}\]

OpenStudy (waheguru):

My question is how we get rid of the parenthesisi

OpenStudy (waheguru):

Do we remove them after we add what ever is inside it

OpenStudy (anonymous):

Yes..

OpenStudy (eyust707):

yes.. if you can add whats inside

OpenStudy (eyust707):

if not you may need to distribute or they could possibly be pointless

OpenStudy (waheguru):

SO In this case -50 - 4 ( (-10-4 (-3+1)^2 ) / (-2) ) oncre the -3 and the 1 are added dont we remove the brackets to make it -50 - 4 ( (-10-4 +2 ^2 ) / (-2) )

OpenStudy (waheguru):

rather than multiplyingg

OpenStudy (anonymous):

What is the inner most bracket in the expression??

OpenStudy (waheguru):

(-3+1) is the inner most bracket

OpenStudy (anonymous):

But it is having an exponent too... so after adding -3 and 1 you cannot eliminate this bracket.. You will get: \[(-2)^2\] Now apply square..

OpenStudy (eyust707):

keep the parenthesis until after you evaluate the exponent. When you do finally remove the parenthesis if there is a number directly next the the parenthesis it becomes multiplication.

OpenStudy (waheguru):

Is that a rule keeping the parentsiis until the exponent has been evaulated

OpenStudy (waheguru):

*parenthesis

OpenStudy (eyust707):

yes!

OpenStudy (eyust707):

have you ever been taught the order of operations?

OpenStudy (anonymous):

In this case after taking the square also you cannot eliminate the brackets because there is a number outside too.. \[-4(-3+1)^2\] \[= -4(-2)^2\] \[= -4(4)\] Now after multiplying you can eliminate the brackets... Like: \[-4(4)\] \[= -16\]

OpenStudy (waheguru):

Yea but not the rule of parenthesis

OpenStudy (zepp):

-50 - 4 ( (-10-4 (-3+1)^2 ) / (-2) ) \[-50-4*(\frac{-10-4(-3+1)^2}{-2})\]

OpenStudy (waheguru):

Thank you so much that was exacly what i was confused about